Created by Main.BorisBellesia on 2006-12-07 - 19:02
Comments:
Advancement of today:
- AB-PO activities on cells C12_R7 - C16-22_R7
- Polarity tests on 120A converter alimenting cells C6-7-8-9-10_R7
- Disconnection of the circuits of the 120A converter alimenting cells C6-7-8-9-10_R7
- MME activities postponed to Monday
Tomorrow activities:
- AB-PO activities on cells C24-30_R7
- 4 hr run in C12_R7-C22_R7 and C32-C34_R7 and C28-C32_L8 simultaneously (shortened version of the 24 hr run)
- ELQA tests - whole day / no co-activities
Edit, attach file or add comment
Created by Main.MircoCoccoli on 2006-12-07 - 12:37

RAT_20061207.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2006x12x07x12x29/RAT_20061207.pdf
Open Issues:
AC non-conformity
- 13.11.06 Non-conformity of the crates of cryo instrumentation (inrush current) (A.Suraci: ready next year)
- 13.11.06 Non-conformity of the AC cabling of the crates under the magnets. This concerns ACR, MEL, orbit corrector power converters & VAC.
DC cabling
- 30.11.06 Non conformity: 4 temperature sensor cables need repair (D.Bozzini, A.Suraci)
- 25.11.06 Functional test on the leads heaters: DFBAO, DFBMA, DFBMC, DFBXG
- 22.11.06 Pre-connection of 120A cables in LSSR7 (TS/EL)
- 20.11.06 Galvanic insulation installation on all DFBs (AT/MEL)
- 13.11.06 6 kA and 13 kA cables positioning at DFBAO, DFBMA, DFBMC (Flohe)
Vacuum
-
Cryogenics
- 22.11.06 Cryo-valves remote control in CCC by AB/CO
- 22.11.06 Control of quench valve
DFB Commissioning
- 04.12.06 Re-positioning collars used to fix the WRL to the CLs (all DFBs) (A.Perin-D.Bozzini-F.Millet)
- 29.11.06 Pumping of Current Leads insulation vacuum for DFBX
- 13.11.06 ElQA-TP4-A for all DFBs between 7 and 8 (D.Bozzini: next(w48): DFBAO HC module, DFBMA, DFBMC, DFBX))
Safety
General remark: written communication in advance to announce operations!!!
- 25.11.06 ElQA-TP4B safety procedure (1900 volts tests!).
- 23.11.06 Wooden structure for UA access restriction (that will allow transport).
- 23.11.06 Preparation of a prototype insulating protection (derived from a plastic tube) of the monorail by J-C.Guillaume and approval by J.Etheridge.
- 23.11.06 Water leak on the tunnel concrete wall to be fixed (C33L8).
- 16.11.06 Cool down safety procedure & access conditions
Preparation of powering test
- 30.11.06 El_QA performed on C16L8. ICC test showed reversed sequence of V-taps on circuit RCBV16.L8B1. To be discussed on next RAT.
- 22.11.06 Installation of flexible cables by CV (UA83)
- 16.11.06 Water circulating in UA83 for the Powering Procedure Test: filters need to be checked.
AOB
- 13.11.06 BPM connection in Q2, Q3, Q4, Q5, Q6, Q7–wk49 (J-C.Guillaume & R.Jones)
MircoCoccoli - 07 Dec 2006 - 12:36
Edit, attach file or add comment
Created by Main.MirkoPojer on 2006-12-07 - 12:36

RAT_20061207.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint5x2006x12x07x12x32/RAT_20061207.pdf
Comments: Next meeting in CCC on Monday Dec,11.
Open Issues
- 22.11 Fire detection – awaiting certification
- 01.12 Short cables on 13 kA EE to be changed later on (temporary solution in place)
- 06.12 Permanent labels for RQ converters in RR57 (according to ECR)
- 07.12 Cables re-distribution for 600 A
- 07.12 Access key to electrical distribution room in UJ56
Heat Runs and Test Schedule
- 07.12 Fire detection test
- 08.12 Full power ventilation tests in RR57 – from 8:30 till 12:00
- 08.12 Afternoon: Polarity Test_1st part
- 11.12 10am: 24-h heat run
- 12.12 UPS test.
- 13.12 Electric cut (AUG simulation).
- 13.12 Polarity Test_2nd part
- 14.12 13kA energy extraction test.
Edit, attach file or add comment